Dash Company adopted a standard costing system several years ago. The standard costs for the prime costs (.e., direct materiaRequired 1 Required 2 1. Compute for November: a. The direct labor efficiency variance. Is this variance favorable (F) or unfRequired 1 Required 2 2. Prepare journal entries to record all transactions, including the variances in requirement 1. (If noJournal entry worksheet Record the entry for the issue of materials to production. Note: Enter debits before credits. TransacRecord the entry for direct labor wages and standard direct labor cost incurred for manufacture. Note: Enter debits before cr
Answer #1

Direct materials

Direct labor

Units completed

7500

7500

Equivalent units in ending WIP Inventory

600

450 (600*75%)

Equivalent units produced, November

8100

7950

Standard quantity of input per unit of output

5

6

Total standard quantity for November production

40500

47700

Part 1 a

Labor efficiency variance = SP ×(AQ −SQ) = 16*(46200-47700)=-24000 = $24000 Favorable (F)

Part 1 B

Labor rate (price) variance = (AP ×AQ) −(SP ×AQ)= $680,000 −($16/hr. ×46200 hrs.)= -59200 = $59200 Favorable (F)

Part 1 C

Actual number of kilograms of materials used in November = standard quantity allowed + per unit materials usage variance (unfavorable) = 40500+(3600/5) = 41220 kilograms

Part 1 D

Materials price variance = total materials variance −materials usage variance = 840 U – 3600 U =$2760 F

Therefore,

Actual price per kilogram of material (AP) = Standard price per kg. (SP) −Per unit favorable price variance = $6.00 per kg. −($2760 ÷69000 kgs.) = $6.00 – 0.04 = $5.96 per kilogram

Part 1 E

Amount of direct materials cost and direct labor cost transferred to the finished goods inventory account = Standard manufacturing cost/unit * number of units manufactured in November = (30+96)*7500 = $945000

Part 1 F

Total amount of direct materials cost and direct labor cost in the work-in-process inventory account at the end of the month = (30*600)+(19.60*450) = $26820

Part 2

No.

General journal

Debit

Credit

A

Materials Inventory (5*69000)

345000

Materials Purchase-Price Variance

2760

Accounts Payable

342240

B

Work-in-Process Inventory (8100*30)

243000

Materials Usage Variance

36900

Materials Inventory (41220*5)

206100

C

Work-in-Process Inventory (7950*96)

763200

Labor Rate Variance

24000

Labor Efficiency Variance

59200

Accrued payroll

680000